# Revisit What You See: Disclose Language Prior in Vision Tokens for Efficient Guided Decoding of LVLMs

### v1

@@ 5.3 Intra-alignment
Improving visual grounding (e.g., mitigating object hallucination) of LVLMs during decoding has evolved through distinct methodological paradigms, each addressing unique aspects of visual-textual misalignment. The research trajectory begins with Contrastive Decoding [38] strategies, pioneered by VCD [23], which introduced input distortions to contrast original and perturbed visual distributions, thereby reducing reliance on language priors within LVLMs. This foundation expanded into instruction-aware variants such as ICD [43], which amplifies alignment uncertainty through textual prompt disturbances, and M3ID [24], which identifies language priors in the LLM decoder via vision-agnostic input perturbations. Concurrently, attention-centric approaches emerged [44, 45, 46, 25, 26], aiming to mitigate attention misallocation by recalibrating cross-modal attention or introducing contrastive signals for token-level grounding. Other paradigms [47, 48, 49] have also been explored in recent literature, but all these intra-alignment approaches require additional computation to obtain vision-related reference information.
6 Conclusion
In this paper, we proposed ReVisiT, a simple yet effective LVLM decoding method that references vision tokens to guide the text generation process. Unlike prior approaches that introduce extra computational cost to obtain the reference logit at every decoding step, ReVisiT leverages the intrinsic vision token representations with negligible overhead. By dynamically constraining the output vocabulary based on output context, ReVisiT enhances visual grounding and mitigates object hallucinations across both generative and discriminative tasks.
Limitations.
Appendix A Detailed Discussions
A.1 Dataset and evaluation setup
CHAIR evaluation.
CHAIR [27] is a generative benchmark designed to measure object hallucinations in image captioning. A hallucinated object refers to any entity mentioned in the generated caption that is not present in the corresponding image. This metric is widely adopted to evaluate the visual grounding capability of vision-language models. We follow the standard CHAIR evaluation procedure based on the object annotations from the MS COCO dataset [39]. Each image is associated with a set of ground-truth objects defined by the 80 category COCO detection vocabulary. During evaluation, noun phrases extracted from generated captions are matched against this predefined object list. Mentions of out-of-vocabulary objects are ignored, while in-vocabulary objects absent from the ground-truth annotations are considered hallucinated. We report two standard metrics provided by the CHAIR benchmark and one additional metric:
$\begin{split}\text{CHAIR}_{S}&=\frac{\left|\{\text{hallucinated captions}\}\right|}{\left|\{\text{total captions}\}\right|},\\ \text{CHAIR}_{I}&=\frac{\left|\{\text{hallucinated object mentions}\}\right|}{\left|\{\text{total object mentions}\}\right|},\\ \text{Recall}&=\frac{\left|\{\text{correctly predicted objects}\}\right|}{\left|\{\text{total ground-truth objects}\}\right|}\end{split}$
We use the official CHAIR evaluation code [^2] https://github.com/LisaAnne/Hallucination and randomly sample 500 images from the MS COCO validation set using a fixed seed (seed=42) for reproducibility. For each image, we prompt the model with "Please describe this image in detail." to generate captions.
POPE evaluation.
POPE [28] is a discriminative benchmark designed to evaluate the object-level visual grounding capabilities of LVLMs. It formulates the task as a binary object presence question-answering problem, where the model is asked to determine whether a specific object is present in a given image. Evaluation is conducted using standard classification metrics derived from the confusion matrix, including accuracy, precision, and F1 score. Accuracy measures the overall proportion of correct predictions, while precision captures the fraction of predicted positive instances that are indeed correct. F1 score, defined as the harmonic mean of precision and recall, provides a balanced assessment of model performance on this binary classification task. The benchmark covers three datasets: MS COCO [39], A-OKVQA [40], and GQA [41]. For each dataset, POPE defines three types of query strategies—random, popular, and adversarial—yielding a total of nine evaluation scenarios. Binary questions are constructed in the form “Is there a <object> in the image?”, with the queried object selected according to the designated strategy. Random queries are uniformly sampled from the dataset’s object vocabulary. Popular queries focus on frequently occurring objects, while adversarial queries target objects that are semantically plausible in the given context but are absent from the image, thus probing the model’s reliance on prior co-occurrence statistics rather than visual evidence. Ground-truth labels indicating object presence are provided by the official POPE benchmark [^3] https://github.com/AoiDragon/POPE. For MS COCO, labels are derived from object detection annotations, whereas for A-OKVQA and GQA, they are obtained via SEEM-based segmentation annotations. Each query type consists of 3,000 binary QA instances, evenly split between positive and negative samples. As a result, we evaluate across a total of nine scenarios comprising 27,000 QA examples.
AMBER evaluation.
AMBER [29] is a comprehensive benchmark designed to assess hallucination phenomena in LVLMs across multiple dimensions. Unlike CHAIR and POPE that focus on object existence, AMBER evaluates three core hallucination types: existence (fabricating non-existent entities), attribute (misrepresenting object properties), and relation (incorrectly describing inter-object interactions). The benchmark supports both generative tasks and discriminative tasks. We focused on generative setting, where AMBER calculates hallucination rates through automated object matching against ground-truth annotations, similar to CHAIR but extended to attribute and relation verification. We report four metrics provided in the AMBER benchmark to quantify hallucination phenomena in the generative setting:
  • CHAIR measures the proportion of hallucinated object mentions in the response. Given a generated object set $R^{\prime}_{\text{obj}}$ and the ground-truth annotated object set $A_{\text{obj}}$, it is defined as:
    $\text{CHAIR}(R)=1-\frac{|R^{\prime}_{\text{obj}}\cap A_{\text{obj}}|}{|R^{\prime}_{\text{obj}}|}.$
  • Cover evaluates the object coverage of the model output, measuring how well the response mentions the ground-truth objects:
    $\text{Cover}(R)=\frac{|R^{\prime}_{\text{obj}}\cap A_{\text{obj}}|}{|A_{\text{obj}}|}.$
  • Hal is a binary indicator of hallucination presence. A response is considered hallucinatory if any hallucinated object is detected:
    $\text{Hal}(R)=\begin{cases}1&\text{if }\text{CHAIR}(R)\neq 0\\ 0&\text{otherwise}\end{cases}.$
  • Cog captures the proportion of hallucinated objects that belong to a predefined cognitively plausible hallucination set $H_{\text{obj}}$:
    $\text{Cog}(R)=\frac{|R^{\prime}_{\text{obj}}\cap H_{\text{obj}}|}{|R^{\prime}_{\text{obj}}|}.$
All metrics are averaged over the full validation split comprising 1,004 generative samples. Lower values of CHAIR and Hal indicate fewer hallucinations, while higher Cover and Cog scores reflect better object recall and alignment with cognitively plausible patterns. We use the official AMBER benchmark [^4] https://github.com/junyangwang0410/AMBER without modification. For Qwen2.5-VL experiments, input images exceeding a resolution of $1400\times 1400$ were resized to fit GPU memory constraints, due to its dynamic high-resolution encoding.
LLaVA-Bench-In-the-Wild dataset.
We additionally employ the LLaVA-Bench-In-the-Wild [6] dataset to qualitatively evaluate the visual grounding ability of LVLMs in open-ended image understanding. The dataset consists of 24 diverse images, including both real-world photographs and abstract illustrations, and 60 textual prompts designed to elicit complex visual reasoning and language generation. Following prior work [23], we present representative qualitative examples to illustrate how different decoding strategies affect grounding fidelity and response quality in challenging scenarios.
@@ A.1 Dataset and evaluation setup
Input: LVLM $\mathcal{M}$ with final decoder layer $L$, projection (language modeling) head $W_{\texttt{proj}}$, and vocabulary $\mathcal{V}$), input image $I$, set of candidate decoder layers $\mathcal{J}$, input text prompt $x_{\text{raw}}$
@@
Construct input sequence: $\mathbf{z}=\mathbf{v}\mathbin{\|}\mathbf{x}$triangleright $\mathbf{z}=\{z_{0},\dots,z_{T-1}\}$ where $T=|\mathbf{v}|+|\mathbf{x}|$
Initialize output sequence $\mathbf{y}\leftarrow\emptyset$triangleright $t\leftarrow 0$
@@ Appendix B Detailed Quantitative Results
B.1 Comprehensive POPE results
Table 6: Comprehensive Results on POPE benchmark. Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance.
Setup Method LLaVA-1.5 Qwen2.5-VL
Acc. $\uparrow$ Prec. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ Prec. $\uparrow$ F1 $\uparrow$
MS-COCO Random Greedy 89.07 89.54 89.00 84.33 99.71 81.47
DoLa 89.00 89.58 88.92 81.40 99.89 77.17
VCD 87.23 86.62 87.34 84.50 98.96 81.81
M3ID 89.07 89.54 89.00 84.27 99.42 81.42
SID 89.73 91.22 89.55 86.00 99.09 83.85
Ours 89.10 89.82 89.00 84.67 99.62 81.95
Popular Greedy 85.63 83.72 86.03 83.77 98.10 80.92
DoLa 85.63 83.85 86.00 81.10 98.95 76.89
@@ B.1 Comprehensive POPE results
M3ID 85.63 83.72 86.03 83.77 98.01 80.94
SID 86.03 84.71 86.29 85.33 97.32 83.21
Ours 85.80 84.16 86.13 84.17 98.21 81.47
Adversarial Greedy 79.27 74.72 81.01 83.43 97.18 80.61
DoLa 79.40 74.97 81.08 80.83 98.03 76.65
VCD 77.67 72.90 79.77 83.50 96.31 80.85
M3ID 79.27 74.72 81.01 83.40 97.00 80.59
SID 80.20 76.22 81.60 84.33 94.86 82.25
Ours 79.90 76.21 81.22 83.70 97.11 80.99
A-OKVQA Random Greedy 86.33 81.00 87.42 87.87 96.86 86.58
DoLa 86.43 81.25 87.47 84.97 97.73 82.65
VCD 83.70 77.91 85.23 88.43 96.91 87.28
M3ID 86.33 81.00 87.42 87.30 96.05 85.97
SID 87.07 82.55 87.91 87.87 96.18 86.67
Ours 87.37 84.75 87.83 88.20 97.36 86.94
Popular Greedy 78.77 71.74 81.72 86.07 92.73 84.89
DoLa 79.00 72.06 81.85 83.07 92.91 80.87
VCD 76.47 69.58 79.99 86.20 92.36 85.12
M3ID 78.77 71.74 81.72 85.80 92.22 84.63
SID 80.33 73.82 82.70 86.33 92.71 85.23
Ours 81.97 77.54 83.31 86.17 92.42 85.07
Adversarial Greedy 68.20 61.86 74.91 81.27 83.26 80.69
DoLa 68.33 62.00 74.95 79.37 84.77 77.63
VCD 67.80 61.69 74.47 81.53 83.12 81.08
M3ID 68.20 61.86 74.91 80.23 81.51 79.82
SID 70.90 64.30 76.36 81.57 83.37 81.06
Ours 72.83 67.00 76.81 81.40 83.31 80.85
GQA Random Greedy 85.97 79.72 87.30 87.50 97.15 86.08
DoLa 86.07 79.92 87.36 83.47 97.36 80.62
VCD 83.77 77.16 85.53 87.87 96.86 86.58
M3ID 86.00 79.74 87.33 87.27 96.82 85.82
SID 85.97 79.72 87.30 86.80 96.94 85.20
Ours 87.13 83.47 87.80 87.83 97.17 86.50
Popular Greedy 73.80 66.38 78.64 84.40 90.12 83.20
DoLa 73.97 66.56 78.73 79.60 87.76 77.13
VCD 71.87 64.76 77.32 84.33 89.07 83.32
M3ID 73.83 66.39 78.67 82.20 85.88 81.24
SID 73.80 66.38 78.64 83.47 89.34 82.13
Ours 78.67 72.42 81.28 84.63 90.37 83.46
Adversarial Greedy 68.17 61.60 75.19 81.80 84.97 80.94
DoLa 68.40 61.80 75.30 78.80 86.00 76.44
VCD 67.63 61.22 74.82 81.93 84.66 81.19
M3ID 68.20 61.62 75.22 81.17 83.95 80.36
SID 68.17 61.60 75.19 81.17 84.76 80.14
Ours 73.40 66.91 77.68 82.17 85.14 81.38
B.2 Token-Level computation speed
@@
Greedy 25.9 ± 0.2 71.7 ± 0.5
DoLa 33.5 ± 0.3 82.2 ± 0.4
VCD 51.6 ± 0.4 143.8 ± 0.9
M3ID 49.5 ± 0.1 143.8 ± 9.8
SID 51.9 ± 0.4 148.4 ± 0.8
Ours 26.5 ± 0.2 73.2 ± 0.4
@@ C.1 Case studies for vision token analysis
To support the qualitative visualization presented in Figures 1 and 2, we provide the detailed numerical values and positional metadata used in our vision token analysis. All examples are drawn from the LLaVA-Bench-In-the-Wild dataset, using the LLaVA-1.5-7B. Vision token embeddings are extracted from decoder layer $L={32}$, and the threshold for constrained vocabulary selection is set to $\alpha=10^{-5}$.
@@
1 three $-1.04$ $35.45$ $-1.47$ $22.98$ $-2.51$ $58.27$
2 four $-0.68$ $50.78$ $-2.25$ $10.52$ $-2.93$ $38.21$
3 two $-4.57$ $1.04$ $-1.19$ $30.44$ $-5.76$ $2.26$
4 five $-2.54$ $7.91$ $-4.45$ $1.16$ $-6.99$ $0.66$
5 a $-4.33$ $1.31$ $-3.08$ $4.58$ $-7.42$ $0.43$
6 six $-3.94$ $1.94$ $-5.15$ $0.58$ $-9.10$ $0.08$
7 $-5.10$ $0.61$ $-4.22$ $1.47$ $-9.32$ $0.06$
8 several $-5.75$ $0.32$ $-5.39$ $0.45$ $-11.15$ $0.01$
9 seven $-5.79$ $0.31$ $-6.12$ $0.22$ $-11.90$ $0.00$
10 un $-8.65$ $0.02$ $-3.80$ $2.24$ $-12.45$ $0.00$
11 at $-8.28$ $0.03$ $-4.77$ $0.85$ $-13.05$ $0.00$
12 one $-10.05$ $0.00$ $-3.69$ $2.50$ $-13.74$ $0.00$
13 in $-11.15$ $0.00$ $-2.99$ $5.01$ $-14.14$ $0.00$
14 eight $-7.38$ $0.06$ $-6.78$ $0.11$ $-14.16$ $0.00$
15 f $-9.52$ $0.01$ $-4.76$ $0.86$ $-14.28$ $0.00$
16 multiple $-7.66$ $0.05$ $-6.66$ $0.13$ $-14.32$ $0.00$
17 only $-8.68$ $0.02$ $-5.89$ $0.28$ $-14.57$ $0.00$
18 nine $-7.71$ $0.04$ $-6.97$ $0.09$ $-14.68$ $0.00$
19 half $-8.62$ $0.02$ $-6.22$ $0.20$ $-14.85$ $0.00$
20 many $-8.70$ $0.02$ $-6.21$ $0.20$ $-14.91$ $0.00$
21 all $-10.38$ $0.00$ $-4.85$ $0.78$ $-15.23$ $0.00$
22 still $-11.11$ $0.00$ $-4.15$ $1.57$ $-15.26$ $0.00$
23 an $-10.33$ $0.00$ $-5.10$ $0.61$ $-15.43$ $0.00$
24 some $-10.05$ $0.00$ $-5.40$ $0.45$ $-15.46$ $0.00$
25 more $-11.15$ $0.00$ $-4.68$ $0.92$ $-15.83$ $0.00$
26 about $-10.05$ $0.00$ $-5.90$ $0.27$ $-15.95$ $0.00$
27 both $-10.74$ $0.00$ $-5.23$ $0.53$ $-15.98$ $0.00$
28 the $-11.98$ $0.00$ $-4.07$ $1.71$ $-16.05$ $0.00$
29 not $-11.58$ $0.00$ $-4.59$ $1.02$ $-16.16$ $0.00$
30 over $-11.99$ $0.00$ $-4.25$ $1.42$ $-16.25$ $0.00$
31 different $-11.93$ $0.00$ $-4.73$ $0.88$ $-16.66$ $0.00$
32 ten $-8.56$ $0.02$ $-8.15$ $0.03$ $-16.71$ $0.00$
33 just $-10.97$ $0.00$ $-5.78$ $0.31$ $-16.75$ $0.00$
34 fruit $-11.81$ $0.00$ $-5.04$ $0.65$ $-16.85$ $0.00$
35 as $-12.06$ $0.00$ $-4.82$ $0.81$ $-16.88$ $0.00$
36 small $-12.11$ $0.00$ $-4.86$ $0.78$ $-16.97$ $0.00$
37 various $-10.39$ $0.00$ $-6.71$ $0.12$ $-17.10$ $0.00$
38 also $-11.94$ $0.00$ $-5.21$ $0.55$ $-17.15$ $0.00$
39 very $-11.80$ $0.00$ $-5.54$ $0.39$ $-17.35$ $0.00$
40 around $-11.28$ $0.00$ $-6.07$ $0.23$ $-17.35$ $0.00$
41 numerous $-10.02$ $0.00$ $-7.51$ $0.05$ $-17.54$ $0.00$
42 no $-10.70$ $0.00$ $-6.93$ $0.10$ $-17.64$ $0.00$
43 total $-11.17$ $0.00$ $-6.80$ $0.11$ $-17.97$ $0.00$
44 currently $-11.17$ $0.00$ $-6.89$ $0.10$ $-18.06$ $0.00$
45 cut $-11.78$ $0.00$ $-6.46$ $0.16$ $-18.24$ $0.00$
46 quite $-11.62$ $0.00$ $-6.71$ $0.12$ $-18.33$ $0.00$
47 few $-11.91$ $0.00$ $-6.64$ $0.13$ $-18.55$ $0.00$
48 twelve $-10.07$ $0.00$ $-8.53$ $0.02$ $-18.60$ $0.00$
49 lots $-11.52$ $0.00$ $-7.37$ $0.06$ $-18.89$ $0.00$
50 th $-11.34$ $0.00$ $-7.63$ $0.05$ $-18.98$ $0.00$
51 actually $-11.84$ $0.00$ $-7.40$ $0.06$ $-19.24$ $0.00$
52 approximately $-11.05$ $0.00$ $-8.45$ $0.02$ $-19.50$ $0.00$
53 exactly $-12.14$ $0.00$ $-7.48$ $0.06$ $-19.62$ $0.00$
54 plenty $-11.78$ $0.00$ $-7.95$ $0.04$ $-19.73$ $0.00$
55 eleven $-10.82$ $0.00$ $-9.25$ $0.01$ $-20.07$ $0.00$
[…]

### v2

@@ A.3.3 Intra-alignment
Improving visual grounding (e.g., mitigating object hallucination) of LVLMs during decoding has evolved through distinct methodological paradigms, each addressing unique aspects of visual-textual misalignment. The research trajectory begins with Contrastive Decoding Li et al. (2023a) strategies, pioneered by VCD Leng et al. (2024), which introduced input distortions to contrast original and perturbed visual distributions, thereby reducing reliance on language priors within LVLMs. This foundation expanded into instruction-aware variants such as ICD Wang et al. (2024), which amplifies alignment uncertainty through textual prompt disturbances, and M3ID Favero et al. (2024), which identifies language priors in the LLM decoder via vision-agnostic input perturbations. Concurrently, attention-centric approaches emerged Gong et al. (2024); Liu et al. (2024b); Woo et al. (2024); Huo et al. (2025); Kang et al. (2025), aiming to mitigate attention misallocation by recalibrating cross-modal attention or introducing contrastive signals for token-level grounding. Other paradigms Chen et al. (2024a); Huang et al. (2024); Wang et al. (2025) have also been explored in recent literature, but all these intra-alignment approaches require additional computation to obtain vision-related reference information.
@@
Input: LVLM $\mathcal{M}$ with final decoder layer $L$, projection (language modeling) head $W_{\texttt{proj}}$, and vocabulary $\mathcal{V}$), input image II, set of candidate decoder layers $\mathcal{J}$, input text prompt $x_{\text{raw}}$
@@
Construct input sequence: $\mathbf{z}=\mathbf{v}\mathbin{\|}\mathbf{x}$ $\triangleright$ $\mathbf{z}=\{z_{0},\dots,z_{T-1}\}$ where $T=|\mathbf{v}|+|\mathbf{x}|$
Initialize output sequence $\mathbf{y}\leftarrow\emptyset$ $\triangleright$ $t\leftarrow 0$
@@ Appendix B Detailed Quantitative Results
B.1 Comprehensive results for 7 models
Table 9: Comprehensive results on CHAIR benchmark across various model sizes and architectures. Lower scores ($\downarrow$) on CHAIRS, CHAIRI and higher ($\uparrow$) F1 indicate better performance. The best results in each model are bolded.
Model Method CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
LLaVA-1.5-7B Greedy 53.8 14.66 82.33
[…]
Qwen2.5-VL-7B Greedy 35.2 8.43 79.85
[…]
InternVL2.5-8B Greedy 33.4 8.47 80.58
[…]
LLaVA-1.5-13B Greedy 48.0 13.29 83.27
[…]
InternVL3-14B Greedy 33.2 9.34 81.49
[…]
InternVL2.5-26B Greedy 31.8 8.39 79.60
[…]
Qwen2.5-VL-32B Greedy 53.6 8.85 82.00
[…]
B.2 Comprehensive Ablation results
Table 10: Ablation study. Panel (a) varies the vision token selection criterion (min/max = min-/max-JSD, random = uniform) and use of a vocabulary subset (full = no constraint, subset = APC). Panel (b) varies the layer scope (last vs all) and the APC threshold $\alpha$. Metrics are CHAIRS (↓), CHAIRI (↓), and F1 (↑). Highlighted rows mark the main configuration.
(a) Vocabulary Constraint $\times$ Token Selection
Vocab Selection CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
subset max 0.0 0.00 0.67
subset random 2.6 7.94 17.63
[…]
full min 0.2 7.14 1.52
(b) Layer Scope $\times$ Threshold $\alpha$
Layer Scope $\bm{\alpha}$ CHAIRS $\downarrow$ CHAIRI $\downarrow$ F1 $\uparrow$
last $1\mathrm{e}{-1}$ 33.8 8.32 80.02
last $1\mathrm{e}{-2}$ 37.0 9.13 79.88
last $1\mathrm{e}{-3}$ 34.0 7.88 80.97
last $1\mathrm{e}{-4}$ 34.2 8.21 80.41
last $1\mathrm{e}{-5}$ 31.4 8.04 80.12
last $1\mathrm{e}{-6}$ 32.8 7.21 79.68
all $1\mathrm{e}{-1}$ 33.8 8.82 80.70
all $1\mathrm{e}{-2}$ 36.8 8.91 79.93
all $1\mathrm{e}{-3}$ 38.8 9.36 79.60
all $1\mathrm{e}{-4}$ 32.8 8.17 80.17
[…]
all $1\mathrm{e}{-6}$ 32.4 9.38 79.90
B.3 Comprehensive POPE results
Table 11: Comprehensive results on POPE benchmark across benchmarks (MS-COCO, A-OKVQA, GQA) and setups (Random / Popular / Adversarial). Higher scores ($\uparrow$) on Accuracy, Precision, and F1 indicate better performance.
Setup Method LLaVA-1.5-7B Qwen2.5-VL-7B
Acc $\uparrow$ Prec $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ Prec $\uparrow$ F1 $\uparrow$
MS-COCO Random Greedy 89.07 89.54 89.00 84.07 99.71 81.09
DoLa 89.00 89.58 88.92 81.53 99.89 77.37
VCD 87.23 86.62 87.34 84.57 99.15 81.88
M3ID 89.07 89.54 89.00 83.93 99.32 80.96
CODE 86.53 87.23 86.41 84.00 98.02 81.26
SID 89.67 91.49 89.43 86.93 99.03 85.10
[…]
MS-COCO Popular Greedy 85.63 83.72 86.03 83.47 97.99 80.52
DoLa 85.63 83.85 86.00 81.10 98.54 76.96
@@ B.3 Comprehensive POPE results
M3ID 85.63 83.72 86.03 83.40 97.81 80.46
CODE 84.73 84.14 84.86 83.63 97.02 80.92
SID 86.13 85.19 86.32 86.23 97.22 84.42
[…]
MS-COCO Adversarial Greedy 79.27 74.72 81.01 83.10 97.06 80.16
DoLa 79.40 74.97 81.08 80.83 97.73 76.71
VCD 77.67 72.90 79.77 83.37 95.96 80.73
M3ID 79.27 74.72 81.01 83.00 96.79 80.06
CODE 78.67 75.23 80.02 83.00 95.41 80.31
SID 80.30 76.56 81.59 85.17 94.74 83.39
[…]
A-OKVQA Random Greedy 86.33 81.00 87.42 87.80 96.86 86.49
DoLa 86.43 81.25 87.47 85.33 96.99 83.26
VCD 83.70 77.91 85.23 87.87 96.56 86.62
M3ID 86.33 81.00 87.42 87.40 95.91 86.11
CODE 83.80 78.48 85.18 87.50 93.64 86.55
SID 87.57 83.48 88.28 88.83 96.12 87.88
[…]
A-OKVQA Popular Greedy 78.77 71.74 81.72 85.93 92.58 84.74
DoLa 79.00 72.06 81.85 83.37 92.17 81.43
VCD 76.47 69.58 79.99 85.90 92.10 84.78
M3ID 78.77 71.74 81.72 85.77 92.21 84.59
CODE 77.70 71.17 80.68 86.07 90.62 85.24
SID 80.63 74.30 82.87 87.00 92.11 86.16
[…]
A-OKVQA Adversarial Greedy 68.20 61.86 74.91 80.83 82.59 80.30
DoLa 68.33 62.00 74.95 79.63 84.22 78.17
VCD 67.80 61.69 74.47 81.10 82.69 80.63
M3ID 68.20 61.86 74.91 79.90 81.00 79.54
CODE 68.30 62.23 74.61 79.37 78.73 79.59
SID 71.33 64.75 76.57 82.40 83.38 82.14
[…]
GQA Random Greedy 85.97 79.72 87.30 87.53 97.15 86.12
DoLa 86.07 79.92 87.36 83.47 94.66 81.10
VCD 83.77 77.16 85.53 87.83 96.78 86.55
M3ID 86.00 79.74 87.33 87.43 96.14 86.12
CODE 84.87 78.58 86.37 87.07 94.48 85.89
SID 85.97 79.72 87.30 88.80 96.78 87.88
[…]
GQA Popular Greedy 73.80 66.38 78.64 83.80 88.82 82.68
DoLa 73.97 66.56 78.73 78.57 83.71 76.80
VCD 71.87 64.76 77.32 83.33 87.09 82.44
M3ID 73.83 66.39 78.67 80.67 82.39 80.14
CODE 73.90 66.60 78.60 81.47 83.29 80.95
SID 73.80 66.38 78.64 84.27 87.25 83.61
[…]
GQA Adversarial Greedy 68.17 61.60 75.19 81.50 84.36 80.70
DoLa 68.40 61.80 75.30 78.37 83.32 76.63
VCD 67.63 61.22 74.82 82.00 84.63 81.29
M3ID 68.20 61.62 75.22 80.50 82.11 80.00
CODE 68.70 62.12 75.39 79.87 80.56 79.64
SID 68.17 61.60 75.19 82.37 83.79 81.99
[…]
B.4 Token-Level computation speed
@@
Greedy 26.0 ± 0.2 72.3 ± 1.5
DoLa 34.2 ± 0.4 82.6 ± 0.4
VCD 51.7 ± 0.4 143.4 ± 1.1
M3ID 49.5 ± 0.2 143.3 ± 13.2
CODE 81.0 ± 6.3 176.2 ± 45.1
SID 51.9 ± 0.4 146.5 ± 1.3
[…]
@@ C.1 Case studies for vision token analysis
To support the qualitative visualization presented in Figures 5 and 2, we provide the detailed numerical values and positional metadata used in our vision token analysis. All examples are drawn from the LLaVA-Bench-In-the-Wild dataset, using the LLaVA-1.5-7B. Vision token embeddings are extracted from decoder layer L=32L={32}, and the threshold for constrained vocabulary selection is set to $\alpha=10^{-5}$.
@@
1 three 1.04-1.04 $35.45$ 1.47-1.47 $22.98$ 2.51-2.51 $58.27$
2 four 0.68-0.68 $50.78$ 2.25-2.25 $10.52$ 2.93-2.93 $38.21$
3 two 4.57-4.57 $1.04$ 1.19-1.19 $30.44$ 5.76-5.76 $2.26$
4 five 2.54-2.54 $7.91$ 4.45-4.45 $1.16$ 6.99-6.99 $0.66$
5 a 4.33-4.33 $1.31$ 3.08-3.08 $4.58$ 7.42-7.42 $0.43$
6 six 3.94-3.94 $1.94$ 5.15-5.15 $0.58$ 9.10-9.10 $0.08$
7 5.10-5.10 $0.61$ 4.22-4.22 $1.47$ 9.32-9.32 $0.06$
8 several 5.75-5.75 $0.32$ 5.39-5.39 $0.45$ 11.15-11.15 $0.01$
9 seven 5.79-5.79 $0.31$ 6.12-6.12 $0.22$ 11.90-11.90 $0.00$
10 un 8.65-8.65 $0.02$ 3.80-3.80 $2.24$ 12.45-12.45 $0.00$
11 at 8.28-8.28 $0.03$ 4.77-4.77 $0.85$ 13.05-13.05 $0.00$
12 one 10.05-10.05 $0.00$ 3.69-3.69 $2.50$ 13.74-13.74 $0.00$
13 in 11.15-11.15 $0.00$ 2.99-2.99 $5.01$ 14.14-14.14 $0.00$
14 eight 7.38-7.38 $0.06$ 6.78-6.78 $0.11$ 14.16-14.16 $0.00$
15 f 9.52-9.52 $0.01$ 4.76-4.76 $0.86$ 14.28-14.28 $0.00$
16 multiple 7.66-7.66 $0.05$ 6.66-6.66 $0.13$ 14.32-14.32 $0.00$
17 only 8.68-8.68 $0.02$ 5.89-5.89 $0.28$ 14.57-14.57 $0.00$
18 nine 7.71-7.71 $0.04$ 6.97-6.97 $0.09$ 14.68-14.68 $0.00$
19 half 8.62-8.62 $0.02$ 6.22-6.22 $0.20$ 14.85-14.85 $0.00$
20 many 8.70-8.70 $0.02$ 6.21-6.21 $0.20$ 14.91-14.91 $0.00$
21 all 10.38-10.38 $0.00$ 4.85-4.85 $0.78$ 15.23-15.23 $0.00$
22 still 11.11-11.11 $0.00$ 4.15-4.15 $1.57$ 15.26-15.26 $0.00$
23 an 10.33-10.33 $0.00$ 5.10-5.10 $0.61$ 15.43-15.43 $0.00$
24 some 10.05-10.05 $0.00$ 5.40-5.40 $0.45$ 15.46-15.46 $0.00$
25 more 11.15-11.15 $0.00$ 4.68-4.68 $0.92$ 15.83-15.83 $0.00$
26 about 10.05-10.05 $0.00$ 5.90-5.90 $0.27$ 15.95-15.95 $0.00$
27 both 10.74-10.74 $0.00$ 5.23-5.23 $0.53$ 15.98-15.98 $0.00$
28 the 11.98-11.98 $0.00$ 4.07-4.07 $1.71$ 16.05-16.05 $0.00$
29 not 11.58-11.58 $0.00$ 4.59-4.59 $1.02$ 16.16-16.16 $0.00$
30 over 11.99-11.99 $0.00$ 4.25-4.25 $1.42$ 16.25-16.25 $0.00$
31 different 11.93-11.93 $0.00$ 4.73-4.73 $0.88$ 16.66-16.66 $0.00$
32 ten 8.56-8.56 $0.02$ 8.15-8.15 $0.03$ 16.71-16.71 $0.00$
33 just 10.97-10.97 $0.00$ 5.78-5.78 $0.31$ 16.75-16.75 $0.00$
34 fruit 11.81-11.81 $0.00$ 5.04-5.04 $0.65$ 16.85-16.85 $0.00$
35 as 12.06-12.06 $0.00$ 4.82-4.82 $0.81$ 16.88-16.88 $0.00$
36 small 12.11-12.11 $0.00$ 4.86-4.86 $0.78$ 16.97-16.97 $0.00$
37 various 10.39-10.39 $0.00$ 6.71-6.71 $0.12$ 17.10-17.10 $0.00$
38 also 11.94-11.94 $0.00$ 5.21-5.21 $0.55$ 17.15-17.15 $0.00$
39 very 11.80-11.80 $0.00$ 5.54-5.54 $0.39$ 17.35-17.35 $0.00$
40 around 11.28-11.28 $0.00$ 6.07-6.07 $0.23$ 17.35-17.35 $0.00$
41 numerous 10.02-10.02 $0.00$ 7.51-7.51 $0.05$ 17.54-17.54 $0.00$
42 no 10.70-10.70 $0.00$ 6.93-6.93 $0.10$ 17.64-17.64 $0.00$
43 total 11.17-11.17 $0.00$ 6.80-6.80 $0.11$ 17.97-17.97 $0.00$
44 currently 11.17-11.17 $0.00$ 6.89-6.89 $0.10$ 18.06-18.06 $0.00$
45 cut 11.78-11.78 $0.00$ 6.46-6.46 $0.16$ 18.24-18.24 $0.00$
46 quite 11.62-11.62 $0.00$ 6.71-6.71 $0.12$ 18.33-18.33 $0.00$
47 few 11.91-11.91 $0.00$ 6.64-6.64 $0.13$ 18.55-18.55 $0.00$
48 twelve 10.07-10.07 $0.00$ 8.53-8.53 $0.02$ 18.60-18.60 $0.00$
49 lots 11.52-11.52 $0.00$ 7.37-7.37 $0.06$ 18.89-18.89 $0.00$
50 th 11.34-11.34 $0.00$ 7.63-7.63 $0.05$ 18.98-18.98 $0.00$
51 actually 11.84-11.84 $0.00$ 7.40-7.40 $0.06$ 19.24-19.24 $0.00$
52 approximately 11.05-11.05 $0.00$ 8.45-8.45 $0.02$ 19.50-19.50 $0.00$
53 exactly 12.14-12.14 $0.00$ 7.48-7.48 $0.06$ 19.62-19.62 $0.00$
54 plenty 11.78-11.78 $0.00$ 7.95-7.95 $0.04$ 19.73-19.73 $0.00$
55 eleven 10.82-10.82 $0.00$ 9.25-9.25 $0.01$ 20.07-20.07 $0.00$